describe a physical situtation in which an object that moves in xy plane might have an acceleration with a constant positve x component and a constant negative y component? 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
For  acceleration with a constant positve x component and a constant negative y component
the position vector must be the function of time with quadratic forrm.
Example suppose the psoiton vector  in x and -ve y dircn. may be:
r = t^2i – t^2j
v = dr/dt = 2ti – 2tj
a = dv/dt = 2i – 2j
Hence it has met the given and desired specifications.